Gas systems project

Europe/Zurich
155/R-047 (CERN)

155/R-047

CERN

10
Show room on map
Beatrice Mandelli (CERN), Frederic Merlet (CERN)

GAS M&O meeting 02 December 2019

Present: Beatrice, Roberto, Fred, Patrick, Kacper, Andrea, LPh, Gianluca, Mara, Federico, Fabio

Indico: https://indico.cern.ch/event/862139/

 

ALICE #2107

  • MTR
    • PLC crashed for the second time after a power-cut: PLC changed with a new one —> Michal is investigating

Reminder:

    • exhaust pressure regulation does not work
    • mixer pressure regulation does not work: Zimmerli not installed
  • CPV
    • Start to reconnect the detectors but in few channels there is overpressure, probably coming from the pipes: LPh is investigating
  • TPC
    • Start to put Neon in the TPC
    • Still leak present
  • TRD

Reminder

    • Purifier: to review if keep the safety valve or the rupture disk
    • Purifier: safety valve at the input where we have Ar or Ar/H2 at high pressure —> LPh to check with safety

 

ATLAS #2102

  • CSC

Reminder:

    • After the power-cut there was not communication with PLC: damage of differential circuit breaker —> to repair
  • MDT
    • Rack 63: again problem on one channel reading: at this point is not the ELMB/motherboard but an electrical problem —> Michal/Patrick to investigare
    • Pump + H2O analysis logic: to discuss with Michal

Reminder:

    • In general: several problems with pressure sensor in MDT. Patrick said to add a resistance in the channel
  • RPC
    • Working on the distribution racks
    • Reminder
    • Mixer instability since 6th November: injection increased of 250 l/h
    • Kacper said nothing seen in the flowcell: you see a step in the total flow of the flowcells on the 05 November —> Kacper to check from where it is going
  • TRT
    • Problem with web-mass —> Michal to check
  • w2 and w3 2020: general power-cut: piquet to stop the system on 05 January

 

CMS #2103

  • RPC
    • ON: RB0, RB+1, RB+2
    • Installation of differential pressure sensors for filter
    • Dummy chambers installation on-going
    • Missing quick manifold and flowcell in barrel +1 and +2
    • PTxx04 was oscillating of ~1 mbar: this was due to the MFC of the exhaust that was flushing —> PID adjusted
    • Bubblers of back-up disconnected now that we are in N2
  • DT
    • Difficult to re-start the gas system after power-cut —> to do some tests

Reminder:

    • Test for power cut and re-start for the exhaust line —> to do other tests
  • CSC
    • Install of differential pressure sensor on the pump and Krohne flowmeter done

Reminder:

    • Warning to set-up —> Beatrice/Michal/Roberto
  • GEM
    • Ga system working fine
    • Rack 61 with detector, rack 62 with all by-pass
    • Flowcells problem in rack 61: ch 3, 5, 7, probably a leak in the connectors of the patch panel of GEM —> Mara to check with Daniel
  • R134a recuperation
    • To understand logic in case of presence of leak/iC4H10 —> Roberto/Nicola Broca

 

LHCb #2108

  • RICH1
    • in Run with CO2
    • 1.5 ppm of O2 every 2 days

Reminder

    • Re-fill with CO2 but we take 600 ppm because the pump took Air —> to think what it can be done to avoid Air intake at the re-start of the system
  • RICH2
    • from 27th re-start recuperation
    • now we are at 2.7% of CF4 left in the system
  • SciFi
    • Commissioning on-going

 

PIQUET

  • General power-cut

 

GENERAL

  • LPh show the new plan of CLOUD for the dismounting
    • Reconstruction in 2020 —> LPh to give to Beatrice the time and dates
  • Overview of planning LS2 for the month of December. To add
    • MTR modification back-up
    • MicroMegas
  • Overview of projects for 2020-2021 by Roberto
    • R134a recuperation plant
    • SHINE-RPC
    • ND280-TPC
  • Purifier drawing in WInCC-OA is wrong —> to prepare a drawing to show next meeting —> Beatrice
There are minutes attached to this event. Show them.